# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@
A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity
@@
${}^{a}$ Department of Physics and Astronomy
The University of Iowa, Iowa City, IA 52242, USA
${}^{b}$ Department of Mathematics
University of California, Santa Barbara, CA 93106-3080
Thomas-Whitehead (TW) gravity is a recently formulated projectively invariant extension of Einstein-Hilbert gravity. Projective geometry was used long ago by Thomas et. al. to succinctly package equivalent paths encoded by the geodesic equation. Projective invariance in gravity has further origins in string theory through a geometric action constructed from the method of coadjoint orbits using the Virasoro Algebra. A projectively invariant connection arises from this construction, a part of which is known as the diffeomorphism field. TW Gravity exploits projective Gauss-Bonnet terms in the action functional to endow the diffeomorphism field with dynamics, while allowing the theory to collapse to General Relativity in the limit that the diffeomorphism field vanishes and the connection becomes Levi-Civita. In the original formulation of TW gravity, the diffeomorphism field is projectively invariant but not tensorial and the connection is projectively invariant but not affine. In this paper we reformulate TW gravity in terms of projectively invariant tensor fields and a projectively invariant covariant derivative, derive field equations respecting these symmetries, and show that the field equations obtained are classically equivalent across formulations. This provides a ’Rosetta Stone’ between this newly constructed covariant and projective invariant formulation of TW gravity and the original formulation that was manifestly projectively invariant, but not covariant.
1 Introduction
In the past few decades Whitehead’s generalization has been revisited [7, 8, 9, 10, 11] and projective geometry has been investigated in regards to the equivalence principle and cosmology [12, 13, 14, 15] and Einstein metrics in general [16, 17]. In making projective invariance manifest, the aforementioned projective direction extends the ${\rm d}$-dimensional base manifold $\mathcal{M}$ to a ${\rm d}+1$-dimensional manifold $\mathcal{N}$ now known as the Thomas cone. Thomas-Whitehead (TW) gravity is a recently formulated model that extends Einstein-Hilbert gravity to the Thomas cone [18, 19, 20]. Additional dynamical fields are introduced to perform this extension. The resulting TW gravity is projectively invariant and reduces to Einstein-Hilbert gravity upon setting the additional dynamical fields to zero.
The roots of TW gravity go deeper still, to string theory and the Virasoro algebra. The notion of the diffeomorphism field as a dynamical field arises from geometric actions built via the method of co-adjoing orbits [24, 25, 26, 27, 28, 29, 30, 31, 32]. In [33], the method of co-adjoint orbits was used to produce the two-dimensional Wess-Zumino-Witten action [34, 35] and the Polyakov 2D quantum gravity action [36, 37]. More specifically, a vector field arises from the Kac-Moody algebra and is interpreted as a Yang-Mills field $A_{a}$, and a rank-two field $D_{ab}$ arises from the Virasoro algebra and is interpreted as the diffeomorphism field of TW gravity [33, 38, 39, 40, 41, 18, 19, 20, 42]. This interpretation becomes clear from the transformation properties of $A_{a}$ and $D_{ab}$ [33, 23, 42]: $A_{a}$ indeed transforms as a Yang-Mills field and $D_{ab}$ transforms as part of a projective connection of Thomas [1, 2].
In its original formulation [18, 19, 20] TW gravity was written in terms of the non-tensorial, but projectively invariant diffeomorphism field ${\mathcal{D}}_{ab}$, the non-affine but projectively invariant connection $\Pi^{c}{}_{ab}$, and the metric $g_{ab}$. The projective Schouten tensor ${\mathcal{P}}_{ab}$ was introduced in lieu of ${\mathcal{D}}_{ab}$, though this fails to be projectively invariant. Furthermore, the non-affine connection $\Pi^{c}{}_{ab}$ can be expanded in terms of the affine connection $\Gamma^{c}{}_{ab}$, though this connection is not by itself projectively invariant. Can we have the best of both worlds: a projectively invariant tensor field and a projectively invariant affine connection? The answer is yes, and the purpose of this paper is to demonstrate precisely how this is done. Expressing TW gravity in the resulting covariant and manifestly projectively invariant form is analogous to writing the Maxwell Lagrangian in a manifestly Lorentz invariant form using $F_{\mu\nu}$ rather than $\vec{E}$ and $\vec{B}$ fields. More to the point, a covariant and manifestly projective invariant formulation of TW gravity is analogous to the formulation of general relativity (GR) itself in terms of tensors and covariant derivatives rather than connections and parital derivatives.
As this new formulation of TW gravity is covariant, the usual tricks of GR become at our disposal: partial integration with respect to the covariant derivative is now available for instance. This was not available in the original formulation of TW gravity, and so deriving the field equations was quite cumbersome [18, 19, 20]. Now, with a covariant formulation, deriving the equations of motion becomes much more straightforward. Since the new formulation is manifestly projectively invariant, this simplifies how the action and field equations are expressed, much like, again, how GR is simpler to write in terms of the Ricci tensor and metric, rather than expanding fully in terms of derivatives of the metric.
This paper is structured as follows. Section 2 is a review of TW gravity. Here the equations of motion are seen to be either manifestly projectively invariant or manifestly tensorial, but not both simultaneously. Section 3 contains the main results of the paper: the revelation of the covariant and manifestly projective invariant formulation of TW gravity. Here we present a covariant and manifestly projectively invariant derivation of the field equations, utilizing the familiar tensorial tricks of Einstein-Hilbert gravity such as integrating by parts with a covariant derivative. The independent fields we vary are the metric $g^{mn}$, the projectively invariant Schouten tensor ${\bar{{\mathcal{P}}}}_{ab}$, and the projectively invariant trace-removed Palatini tensor ${\tilde{C}}^{c}{}_{ab}$. The tensor ${\tilde{C}}^{c}{}_{ab}$ is the difference between the projectively invariant affine connection $\bar{\Gamma}^{c}{}_{ab}$ and the Levi-Civita connection $\hat{\Gamma}^{c}{}_{ab}$. We demonstrate the on-shell equivalence of the equations of motion across the different formulations in Section 4.
2 Review of TW Gravity
$\Gamma^{a}{}_{bc}\to\Gamma^{a}{}_{bc}+\delta^{a}{}_{(b}v_{c)}$ (2.1)
with $v_{b}$ an arbitrary one form.
2.1 TW Connections
As originally presented in [18, 20], the TW connection $\tilde{\Gamma}^{\alpha}_{\ \beta\gamma}$ on the ${\rm d}+1$-dimensional Thomas cone $\mathcal{N}$ is expressed in terms of the fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ on the ${\rm d}$-dimensional base manifold $\mathcal{M}$
@@ 2.1 TW Connections
where
${\Pi}^{a}_{\,\,\,\,bc}=$ ${\Gamma}^{a}_{\,\,\,\,bc}+\delta^{a}{}_{(c}~{}\alpha_{b)}~{}~{}~{},~{}~{}~{}\alpha_{b}=-\tfrac{1}{{\rm d}+1}\Gamma^{c}{}_{bc}$ (2.3)
$\Upsilon^{\alpha}=$ $(0,0,\dots,0,\lambda)~{}~{}~{},~{}~{}~{}\alpha_{\beta}=\left(\alpha_{c},\lambda^{-1}\right)~{}~{}~{}.$ (2.4)
$x^{\alpha}=$ $(x^{a},\lambda)~{}~{}~{}.$ (2.5)
${\Gamma}^{\prime a}{}_{mn}=$ $\frac{\partial x^{\prime a}}{\partial x^{b}}\frac{\partial x^{p}}{\partial x^{\prime m}}\frac{\partial x^{q}}{\partial x^{\prime n}}{\Gamma}^{b}{}_{pq}+\frac{\partial^{2}x^{b}}{\partial x^{\prime m}\partial^{\prime n}}\frac{\partial x^{\prime a}}{\partial x^{b}}~{}~{}~{}.$ (2.6)
$\alpha^{\prime}_{b}=$ $\frac{\partial x^{m}}{\partial x^{\prime b}}\alpha_{m}-\frac{1}{{\rm d}+1}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\frac{\partial x^{\prime a}}{\partial x^{m}}~{}~{}~{}.$ (2.7)
[…]
$x^{\prime\alpha}=$ $(x^{\prime 0}(x^{m}),x^{\prime 1}(x^{m}),\dots,x^{\prime{\rm d}-1}(x^{m}),\lambda^{\prime}=\lambda J^{1/({\rm d}+1)})~{}~{}~{},~{}~{}~{}J=|\partial x^{m}/\partial x^{\prime n}|$ (2.8)
${\tilde{\Gamma}}^{\prime\alpha}{}_{\mu\nu}=$ $\frac{\partial x^{\prime\alpha}}{\partial x^{\rho}}\frac{\partial x^{\sigma}}{\partial x^{\prime\mu}}\frac{\partial x^{\beta}}{\partial x^{\prime\nu}}{\tilde{\Gamma}}^{\rho}{}_{\sigma\beta}+\frac{\partial^{2}x^{\beta}}{\partial x^{\prime\mu}\partial^{\prime\nu}}\frac{\partial x^{\prime\alpha}}{\partial x^{\beta}}~{}~{}~{}.$ (2.9)
[…]
$\alpha^{\prime}_{\beta}=$ $\frac{\partial x^{\mu}}{\partial x^{\prime\beta}}\alpha_{\mu}=-\frac{1}{{\rm d}+1}\delta^{b}{}_{\beta}\left(\frac{\partial x^{m}}{\partial x^{\prime b}}\Gamma^{n}{}_{mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\right)+\delta^{\lambda}{}_{\beta}\lambda^{\prime-1}~{}~{}~{}.$ (2.10)
This is consistent with $\alpha_{b}$ transforming as the normalized trace of an affine connection on the manifold $\mathcal{M}$, Eq. (2.7), as demonstrated by the expansion following the second equality.
The fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ have the advantage of being projectively invariant under a projective transformation as defined in Eq. (2.1). The drawback of $\mathcal{D}_{ab}$ is that it is not a tensor and the drawback of $\Pi^{a}{}_{bc}$ is that it is not an affine connection. Instead, these transform under general coordinate transformations on $\mathcal{M}$ as
$\mathcal{D}^{\prime}_{ab}=$ $\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\mathcal{D}_{mn}-\partial_{m}j_{n}-j_{m}j_{n}+j_{c}\Pi^{c}{}_{mn}\right]\;,$ (2.11)
${\Pi^{\prime}}^{a}_{\ bc}=$ $\frac{\partial x^{\prime a}}{\partial x^{p}}\frac{\partial x^{m}}{\partial x^{\prime b}}\frac{\partial x^{n}}{\partial x^{\prime c}}\Pi^{p}_{\ mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime b}\ \partial x^{\prime c}}-\delta^{a}_{\ (b}\frac{\partial x^{m}}{\partial x^{\prime c)}}j_{m}$ (2.12)
$j_{m}\equiv$ $\tfrac{1}{{\rm d}+1}\partial_{m}\log J=\tfrac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{m}\partial x^{\prime q}}$ (2.13)
@@
In particular, this means that the resulting gravitational model written in terms of $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ known as Thomas-Whitehead gravity [20] though manifestly projectively invariant does not have field equations that are manifestly tensorial, or even tensorial at all in the case of many of the field equations. The main goal of this paper is to express Thomas-Whitehead gravity in terms of a set of fields that are both manifestly tensorial and manifestly projectively invariant on the manifold $\mathcal{M}$. This will lead to field equations that are likewise manifestly tensorial and manifestly projectively invariant on the manifold $\mathcal{M}$, while at the same time being dynamically equivalent to the original set of field equations derived in [20] at least classically.
@@
~$\tilde{\nabla}_{\alpha}v_{\beta}=$ $\partial_{\alpha}v_{\beta}-\tilde{\Gamma}^{\mu}{}_{\alpha\beta}v_{\mu}~{}~{}~{},~{}~{}~{}\tilde{\nabla}_{\alpha}v^{\beta}=\partial_{\alpha}v_{\mu}+\tilde{\Gamma}^{\beta}{}_{\alpha\mu}v^{\mu}$ (2.14)
$\nabla_{a}v_{b}=$ $\partial_{a}v_{b}-\Gamma^{m}{}_{ab}v_{m}~{}~{}~{},~{}~{}~{}\nabla_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Gamma^{b}{}_{am}v^{m}$ (2.15)
˘$\breve{\nabla}_{a}v_{b}=$ $\partial_{a}v_{b}-\Pi^{m}{}_{ab}v_{m}~{}~{}~{},~{}~{}~{}\breve{\nabla}_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Pi^{b}{}_{am}v^{m}$ (2.16)
@@
$\mathcal{P}_{ab}$ $=\mathcal{D}_{ab}-\partial_{a}\alpha_{b}+\Gamma^{c}_{\ ab}\alpha_{c}+\alpha_{a}\alpha_{b}~{}~{}~{}.$ (2.17)

### v3

# Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

@@
Abstract
Thomas-Whitehead (TW) gravity is a recently formulated projectively invariant extension of Einstein-Hilbert gravity. Projective geometry was used long ago by Thomas et. al. to succinctly package equivalent paths encoded by the geodesic equation. Projective invariance in gravity has further origins in string theory through a geometric action constructed from the method of coadjoint orbits using the Virasoro Algebra. A projectively invariant connection arises from this construction, a part of which is known as the diffeomorphism field. TW gravity exploits projective Gauss-Bonnet terms in the action functional to endow the diffeomorphism field with dynamics, while allowing the theory to collapse to general relativity in the limit that the diffeomorphism field vanishes and the connection becomes Levi-Civita. In the original formulation of TW gravity, the diffeomorphism field is projectively invariant but not tensorial and the connection is projectively invariant but not affine. In this paper we reformulate TW gravity in terms of projectively invariant tensor fields and a projectively invariant covariant derivative, derive field equations respecting these symmetries, and show that the field equations obtained are classically equivalent across formulations. This provides a ’Rosetta Stone’ between this newly constructed covariant and projective invariant formulation of TW gravity and the original formulation that was manifestly projective invariant, but not covariant.
I Introduction
In the past few decades Whitehead’s generalization has been revisited [7, 8, 9, 10, 11] and projective geometry has been investigated in regards to the equivalence principle and cosmology [12, 13, 14, 15] and Einstein metrics in general [16, 17]. In making projective invariance manifest, the aforementioned projective direction extends the ${\rm d}$-dimensional base manifold $\mathcal{M}$ to a ${\rm d}+1$-dimensional manifold $\mathcal{N}$ now known as the Thomas cone. Thomas-Whitehead (TW) gravity is a recently formulated model that extends Einstein-Hilbert gravity to the Thomas cone [18, 19, 20, 21]. Additional dynamical fields are introduced to perform this extension. The resulting TW gravity is projectively invariant and reduces to Einstein-Hilbert gravity upon setting the additional dynamical fields to zero.
The roots of TW gravity go deeper still, to string theory and the Virasoro algebra. The notion of the diffeomorphism field as a dynamical field arises from geometric actions built via the method of coadjoint orbits [25, 26, 27, 28, 29, 30, 31, 32, 33]. In [34], the method of coadjoint orbits was used to produce the two-dimensional Wess-Zumino-Witten action [35, 36] and the Polyakov 2D quantum gravity action [37, 38]. More specifically, a vector field arises from the Kac-Moody algebra and is interpreted as a Yang-Mills field $A_{a}$, and a rank-two field $D_{ab}$ arises from the Virasoro algebra and is interpreted as the diffeomorphism field of TW gravity [34, 39, 40, 41, 42, 18, 19, 20, 21, 43]. This interpretation becomes clear from the transformation properties of $A_{a}$ and $D_{ab}$ [34, 24, 43]: $A_{a}$ indeed transforms as a Yang-Mills field and $D_{ab}$ transforms as part of a projective connection of Thomas [1, 2].
In its original formulation [18, 19, 20, 21] TW gravity was written in terms of the non-tensorial, but projectively invariant diffeomorphism field ${\mathcal{D}}_{ab}$, the non-affine but projectively invariant connection $\Pi^{c}{}_{ab}$, and the metric $g_{ab}$. The projective Schouten tensor ${\mathcal{P}}_{ab}$ was introduced in lieu of ${\mathcal{D}}_{ab}$, though this fails to be projectively invariant. Furthermore, the non-affine connection $\Pi^{c}{}_{ab}$ can be expanded in terms of the affine connection $\Gamma^{c}{}_{ab}$, though this connection is not by itself projectively invariant. Can we have the best of both worlds: a projectively invariant tensor field and a projectively invariant affine connection? The answer is yes, and the purpose of this paper is to demonstrate precisely how this is done. Expressing TW gravity in the resulting covariant and manifestly projective invariant form is analogous to writing the Maxwell Lagrangian in a manifestly Lorentz invariant form using $F_{\mu\nu}$ rather than $\vec{E}$ and $\vec{B}$ fields. More to the point, a covariant and manifestly projective invariant formulation of TW gravity is analogous to the formulation of general relativity (GR) itself in terms of tensors and covariant derivatives rather than connections and partial derivatives.
As this new formulation of TW gravity is covariant, the usual tricks of GR become at our disposal: partial integration with respect to the covariant derivative is now available for instance. This was not available in the original formulation of TW gravity, and so deriving the field equations was quite cumbersome [20, 21]. Now, with a covariant formulation, deriving the equations of motion becomes much more straightforward. Since the new formulation is manifestly projective invariant, this simplifies how the action and field equations are expressed, much like, again, how GR is simpler to write in terms of the Ricci tensor and metric, rather than expanding fully in terms of derivatives of the metric.
This paper is structured as follows. Section II is a review of TW gravity. Here the equations of motion are seen to be either manifestly projective invariant or covariant, but not both simultaneously. Section III contains the main results of the paper: the revelation of the covariant and manifestly projective invariant formulation of TW gravity. Here we present a covariant and manifestly projective invariant derivation of the field equations, utilizing the familiar tensorial tricks of Einstein-Hilbert gravity such as integrating by parts with a covariant derivative. The independent fields we vary are the metric $g^{mn}$, the projectively invariant Schouten tensor ${\bar{{\mathcal{P}}}}_{ab}$, and the projectively invariant trace-removed Palatini tensor ${\tilde{C}}^{c}{}_{ab}$. The tensor ${\tilde{C}}^{c}{}_{ab}$ is the difference between the projectively invariant affine connection $\bar{\Gamma}^{c}{}_{ab}$ and the Levi-Civita connection $\hat{\Gamma}^{c}{}_{ab}$. We demonstrate the on-shell (with equations of motion enforced) equivalence of the equations of motion across the different formulations, summarize the equations of motion, and discuss solutions in Section IV.
II Review of TW Gravity
$\Gamma^{a}{}_{bc}\to\Gamma^{a}{}_{bc}+\delta^{a}{}_{(b}v_{c)}$ (1)
with $v_{b}$ an arbitrary one form. Here and throughout the paper, parenthesis denote symmetrization of enclosed indices as described in Eq. (260) of Appendix A.
II.1 TW Connections
As originally presented in [18, 19, 20, 21], the TW connection $\tilde{\Gamma}^{\alpha}_{\ \beta\gamma}$ on the ${\rm d}+1$-dimensional Thomas cone $\mathcal{N}$ is projectively invariant with respect to projective transformations as in Eq. (1). The TW connection is expressed in terms of the fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ on the ${\rm d}$-dimensional base manifold [^1] Here we are using the labels $\mathcal{M}$ and $\mathcal{N}$ as in [19, 20]. Since the Thomas cone is a bundle over $\mathcal{M}$, the works [11, 18, 21, 44] refer to these manifolds as $M$ and $VM$. $\mathcal{M}$
$\tilde{\Gamma}^{\alpha}{}_{\beta\gamma}=\begin{cases}\tilde{\Gamma}^{a}_{\ bc}=\tensor{\Pi}{{}^{a}_{bc}}\\ \tensor{\tilde{\Gamma}}{{}^{\lambda}_{bc}}=\Upsilon^{\lambda}\mathcal{D}_{bc}\\ \tensor{\tilde{\Gamma}}{{}^{a}_{b\lambda}}=\tensor{\tilde{\Gamma}}{{}^{a}_{\lambda b}}=\alpha_{\lambda}\tensor{\delta}{{}^{a}_{b}}\\ \tensor{\tilde{\Gamma}}{{}^{\lambda}_{b\lambda}}=\tensor{\tilde{\Gamma}}{{}^{\lambda}_{\lambda b}}=\tensor{\tilde{\Gamma}}{{}^{\lambda}_{\lambda\lambda}}=0\\ \end{cases}\,$ (2)
where
${\Pi}^{a}_{\,\,\,\,bc}=$ ${\Gamma}^{a}_{\,\,\,\,bc}+\delta^{a}{}_{(c}~{}\alpha_{b)}~{}~{}~{},~{}~{}~{}\alpha_{b}=-\tfrac{1}{{\rm d}+1}\Gamma^{c}{}_{bc}$ (3)
$\Upsilon^{\alpha}=$ $(0,0,\dots,0,\lambda)~{}~{}~{},~{}~{}~{}\alpha_{\beta}=\left(\alpha_{c},\lambda^{-1}\right)~{}.$ (4)
$x^{\alpha}=$ $(x^{a},\lambda)~{}.$ (5)
We note here that the connection is torsion-free, so symmetric in its lower two indices, and written in the conventions of Roberts [8, 11, 18, 19, 20, 21]. A more general treatment, including a revival of Kaluza-Klein-like solutions, can be found in the recent work [44].
${\Gamma}^{\prime a}{}_{mn}=$ $\frac{\partial x^{\prime a}}{\partial x^{b}}\frac{\partial x^{p}}{\partial x^{\prime m}}\frac{\partial x^{q}}{\partial x^{\prime n}}{\Gamma}^{b}{}_{pq}+\frac{\partial^{2}x^{b}}{\partial x^{\prime m}\partial^{\prime n}}\frac{\partial x^{\prime a}}{\partial x^{b}}~{}.$ (6)
$\alpha^{\prime}_{b}=$ $\frac{\partial x^{m}}{\partial x^{\prime b}}\alpha_{m}-\frac{1}{{\rm d}+1}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\frac{\partial x^{\prime a}}{\partial x^{m}}~{}.$ (7)
@@ II.1 TW Connections
$x^{\prime\alpha}=$ $(x^{\prime 0}(x^{m}),x^{\prime 1}(x^{m}),\dots,x^{\prime{\rm d}-1}(x^{m}),\lambda^{\prime}=\lambda J^{\tfrac{1}{{\rm d}+1}})$ (8)
$J\equiv$ $|\partial x^{m}/\partial x^{\prime n}|$ (9)
${\tilde{\Gamma}}^{\prime\alpha}{}_{\mu\nu}=$ $\frac{\partial x^{\prime\alpha}}{\partial x^{\rho}}\frac{\partial x^{\sigma}}{\partial x^{\prime\mu}}\frac{\partial x^{\beta}}{\partial x^{\prime\nu}}{\tilde{\Gamma}}^{\rho}{}_{\sigma\beta}+\frac{\partial^{2}x^{\beta}}{\partial x^{\prime\mu}\partial^{\prime\nu}}\frac{\partial x^{\prime\alpha}}{\partial x^{\beta}}~{}.$ (10)
@@
$\alpha^{\prime}_{\beta}=\frac{\partial x^{\mu}}{\partial x^{\prime\beta}}\alpha_{\mu}$ (11)
$=\frac{-1}{{\rm d}+1}\delta^{b}{}_{\beta}\left[\frac{\partial x^{m}}{\partial x^{\prime b}}\Gamma^{n}{}_{mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\right]+\delta^{\lambda}{}_{\beta}\frac{1}{\lambda^{\prime}}.~{}~{}~{}~{}~{}$ (12)
This is consistent with $\alpha_{b}$ transforming as the normalized trace of an affine connection on the manifold $\mathcal{M}$, Eq. (7), as demonstrated by the expansion following the second equality.
The fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ have the advantage of being projectively invariant under a projective transformation as defined in Eq. (1). The drawback of $\mathcal{D}_{ab}$ is that it is not a tensor and the drawback of $\Pi^{a}{}_{bc}$ is that it is not an affine connection. Instead, these transform under general coordinate transformations on $\mathcal{M}$ as
$\mathcal{D}^{\prime}_{ab}=$ $\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\mathcal{D}_{mn}-\partial_{m}j_{n}-j_{m}j_{n}+j_{c}\Pi^{c}{}_{mn}\right]$ (13)
${\Pi^{\prime}}^{a}_{\ bc}=$ $\frac{\partial x^{\prime a}}{\partial x^{p}}\frac{\partial x^{m}}{\partial x^{\prime b}}\frac{\partial x^{n}}{\partial x^{\prime c}}\Pi^{p}_{\ mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime b}\ \partial x^{\prime c}}$ (14)
$-\delta^{a}_{\ (b}\frac{\partial x^{m}}{\partial x^{\prime c)}}j_{m}$ (15)
$j_{m}\equiv$ $\frac{1}{{\rm d}+1}\partial_{m}\log J=\frac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{m}\partial x^{\prime q}}$ (16)
@@
In particular, this means that the resulting gravitational model written in terms of $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ known as Thomas-Whitehead gravity [18, 19, 20, 21] though manifestly projective invariant does not have field equations that are covariant, or even tensorial at all in the case of many of the field equations. The main goal of this paper is to express Thomas-Whitehead gravity in terms of a set of fields that are both covariant and manifestly projective invariant on the manifold $\mathcal{M}$. This will lead to field equations that are likewise covariant and manifestly projective invariant on the manifold $\mathcal{M}$, while at the same time being dynamically equivalent to the original set of field equations derived in [20, 21] at least classically.
@@
~$\tilde{\nabla}_{\alpha}v_{\beta}=$ $\partial_{\alpha}v_{\beta}-\tilde{\Gamma}^{\mu}{}_{\alpha\beta}v_{\mu}~{},~{}\tilde{\nabla}_{\alpha}v^{\beta}=\partial_{\alpha}v_{\mu}+\tilde{\Gamma}^{\beta}{}_{\alpha\mu}v^{\mu}$ (17)
$\nabla_{a}v_{b}=$ $\partial_{a}v_{b}-\Gamma^{m}{}_{ab}v_{m}~{},~{}\nabla_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Gamma^{b}{}_{am}v^{m}$ (18)
˘$\breve{\nabla}_{a}v_{b}=$ $\partial_{a}v_{b}-\Pi^{m}{}_{ab}v_{m}~{},~{}\breve{\nabla}_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Pi^{b}{}_{am}v^{m}$ (19)
@@
$\mathcal{P}_{ab}$ $=\mathcal{D}_{ab}-\partial_{a}\alpha_{b}+\Gamma^{c}_{\ ab}\alpha_{c}+\alpha_{a}\alpha_{b}~{}.$ (20)